= 2P(Z > 1 = 2P(Z > 2.95 P -value 2P(Z> 2.95) = 2[1-P(Z < 2.95)]

How to write this formula in Excel?
Answer #1

Ans:

Use normsdist Excel function:

Two tailed p-value:

p-value=2*P(Z>2.95)

or

=2*(1-P(Z<2.95))

=2*(1-normsdist(2.95))

=0.0032